OpenStudy (anonymous):

use the squeeze thermo to determine ; lim x- infinity 3 - sin ( e x) / spuare x ^ 2 + 2

OpenStudy (anonymous):

Is your limit \[ \lim_{x\to \infty} \frac { 3 -\sin( e^x)}{\sqrt{x^2+2}} \]

OpenStudy (anonymous):

\[ 0\le \left | \frac { 3 -\sin( e^x)}{\sqrt{x^2+2}} \right|\le \left | \frac { 3}{\sqrt{x^2+2}} \right| + \left | \frac { 1}{\sqrt{x^2+2}} \right| \] Can you argue to finish it?

OpenStudy (anonymous):

The RHS goes to zero so the quantity in between goes to zero by the squeeze theroem

OpenStudy (anonymous):

If the absolute value of a quantity goes to zero, the quantity itself does the same.

OpenStudy (anonymous):

thanks a lot i understand it must i show througher calculations ?

OpenStudy (anonymous):

sorry i dont understand the squeezze theroem it counts 15 marks in my assignment

OpenStudy (anonymous):

The squeeze theorem says if \[ f(x)\le g(x) \le h(x)\\ if \\ \lim_{x\to a} f(x)=\lim_{x\to a}h(x)=L\\then\\ \lim_{x\to a} g(x)=L \]